OpenStudy (anonymous):

How many permutations can be made of the letters of the word "radar" when taken all at a time? 30 24

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

30?

OpenStudy (calculusfunctions):

Permutations (arrangements) of n objects where p of one type are alike, q of another type are alike, r of another type are alike and so on is given by\[\frac{ n! }{ p!q!r!... }\]

OpenStudy (calculusfunctions):

Example: If we consider the # of permutations of the word TORONTO taken all at a time, then we notice there are 7 letters of which there are 2 T's and 3 O's. Thus the total # of permutations possible are\[\frac{ 7! }{ 2!3! }=210 \]

OpenStudy (calculusfunctions):

CORRECTION:\[\frac{ 7! }{ 2!3! }=420\]

OpenStudy (calculusfunctions):

Yes @hhelpplzzzz your answer of 30 is correct because RADAR has 5 letters of which there are 2 A's and 2 R's. Thus\[\frac{ 5! }{ 2!2! }=30\]